OpenStudy (anonymous):

Find the exact value of the following. Do NOT use labels in your answer and do NOT round your answer. -3(1.2)sqrd=

OpenStudy (anonymous):

@FriedRice do you know?

OpenStudy (anonymous):

confused and im thinking very hard

OpenStudy (anonymous):

is the equation suppose to be \[\sqrt{-3(1.2)}\]

OpenStudy (anonymous):

or\[\sqrt[3]{1.2}\]

OpenStudy (anonymous):

\[\sqrt[-3]{1.2}\]

OpenStudy (anonymous):

pick 1 2 or 3

OpenStudy (anonymous):

1 or 3?